# Sparse-Tuning: Adapting Vision Transformers with Efficient Fine-tuning and Inference

### v1

@@ Abstract
Parameter-efficient fine-tuning (PEFT) has emerged as a popular approach for adapting pre-trained Vision Transformer (ViT) models to downstream applications. While current PEFT methods achieve parameter efficiency, they overlook GPU memory and time efficiency during both fine-tuning and inference, due to the repeated computation of redundant tokens in the ViT architecture. This falls short of practical requirements for downstream task adaptation. In this paper, we propose Sparse-Tuning, a novel tuning paradigm that substantially enhances both fine-tuning and inference efficiency for pre-trained ViT models. Sparse-Tuning efficiently fine-tunes the pre-trained ViT by sparsely preserving the informative tokens and merging redundant ones, enabling the ViT to focus on the foreground while reducing computational costs on background regions in the images. To accurately distinguish informative tokens from uninformative ones, we introduce a tailored Dense Adapter, which establishes dense connections across different encoder layers in the ViT, thereby enhancing the representational capacity and quality of token sparsification. Empirical results on VTAB-1K, three complete image datasets, and two complete video datasets demonstrate that Sparse-Tuning reduces the GFLOPs to 62%-70% of the original ViT-B while achieving state-of-the-art performance. Source code is available at https://github.com/liuting20/Sparse-Tuning.
@@ 1 Introduction
Large-scale Vision Transformer (ViT) models [10, 41, 40, 48, 30] have demonstrated strong generalization capabilities across a wide range of downstream vision tasks. The prevalent approach to adapt these models for specific tasks follows the pretrain-then-finetune paradigm, where models are initially pre-trained on large-scale datasets [8, 38, 32] and then fine-tuned for each downstream task. However, as these pre-trained ViT models continue to scale up [57, 7], fully fine-tuning them becomes more computationally intensive. Additionally, there are risks of catastrophic forgetting and overfitting when fine-tuning on limited downstream datasets [59, 24, 9].
Recently, various parameter-efficient fine-tuning (PEFT) methods [21, 34, 22, 25, 5] have been proposed to address the high computational costs and risks associated with fully fine-tuning large models. By updating additional parameters inserted into the model [21, 5] or appended to the input data [34, 25], PEFT methods can achieve similar or even better performance compared to full fine-tuning. Current PEFT methods [29, 26, 12, 44] focus primarily on efficiently fine-tuning a pre-trained ViT for downstream visual tasks. However, despite achieving fine-tuning efficiency, they fall short in inference efficiency due to the repeated computation of redundant tokens in the Multi-Head Attention (MHA) and Feed-Forward Network (FFN) blocks, failing to meet practical requirements.
Inspired by the model acceleration methods for ViT [49, 51, 36, 37], we aim to reduce redundant tokens to enhance efficiency during both the fine-tuning and inference stages. Since most acceleration methods progressively prune tokens across encoder layers in ViT, we consider establishing interactions between shallow encoder layers, which capture local information, and deep encoder layers, which capture global information. By supplementing global information with local details, we aim to accurately discriminate between informative and uninformative tokens. To this end, in this paper, we introduce Sparse-Tuning, a novel tuning paradigm that achieves efficient fine-tuning and inference for ViT adaptation. Sparse-Tuning efficiently fine-tunes the pre-trained ViT by sparsely adapting the tokens to enable the model to focus on informative regions, thereby enhancing both computational and memory efficiency during the fine-tuning and inference stages. Specifically, we design a novel parameter-free Token Sparsification method that allows the pre-trained ViT to progressively preserve informative tokens while integrating uninformative ones into a representative token, thereby reducing redundant computational costs. Moreover, to mitigate information loss from Token Sparsification and efficiently fine-tune the pre-trained ViT, we propose Dense Adapters, which take multiple features from different encoder layers as inputs to establish dense connections between multiple Token Sparsification steps. With these non-trivial designs, Sparse-Tuning enhances performance while reducing training and inference time, GPU memory consumption, and computational cost for ViT adaptation, compared to most PEFT methods [21, 22, 5, 25, 60], as shown in Figure 1.
@@
To fully evaluate the generalization capability, we conduct extensive experiments on the common PEFT benchmark VTAB-1K [58], three complete image datasets: CIFAR-100 [33], SVHN [13], and Food-101 [2], as well as two complete video datasets: Kinetics-400 (K400) [4] and Something-Something V2 (SSv2) [14]. Empirical results on VTAB-1K demonstrate that with only 11.65 GFLOPs, approximately 66% of the computational cost of the original ViT-B, Sparse-Tuning outperforms all state-of-the-art methods in performance, fine-tuning and inference efficiency. Moreover, Sparse-Tuning achieves superior performance in both image and video recognition on complete datasets while significantly improving both fine-tuning and inference efficiency.
@@ 2.1 Parameter-efficient Fine-tuning
With the trend of scaling up Vision Transformers (ViT) [10, 18, 57, 7] for enhanced performance and generalization, adapting the entire model to downstream tasks becomes increasingly computationally expensive. To mitigate this, parameter-efficient fine-tuning (PEFT) [21, 34, 22, 25, 55, 54] emerges as a strategic approach. PEFT methods updates only a small subset of additional parameters while keeping the majority of the pre-trained model frozen, thereby mitigating the risks of catastrophic forgetting and overfitting. Most PEFT methods methods designed for Transformer [50] can be classified into three types: (1) Partially Tuning [31, 56, 11, 17], that only updates a small subset of inherent parameters while freezing most original parameters. (2) Prompt Tuning [34, 25, 24, 53], that integrates a fixed length of learnable tokens (i.e., prompts) appended with the input data. It only updates the prompts during fine-tuning. (3) Adapter Tuning [21, 5, 52, 39], that only updates additional parameters in the module inserted into the model (i.e., Adapter) during fine-tuning.
@@
While most PEFT methods improve parameter efficiency during the fine-tuning stage of ViT, they often introduce new parameters that hinder inference efficiency. Reparameterization methods, such as LoRA [22] and FacT [29], introduce learnable parameters that can be integrated into the original model through reparameterization during inference. Therefore, these methods can maintain the original model’s inference efficiency. However, current PEFT methods fail to enhance inference efficiency while achieving parameter-efficient fine-tuning, which does not meet the practical needs for adapting the large-size ViT (e.g., ViT-L [10]). In this paper, we aim to enhance the efficiency of both the fine-tuning and inference processes for the pre-trained ViT.
2.2 Model Acceleration for ViT
@@
Recently, numerous works have explored to accelerate the inference efficiency of ViT [15, 49, 36, 6, 37], with most of them aiming to reduce the token redundancy in ViT to decrease computational complexity. For instance, DynamicViT [49] efficiently sparsifies ViT by pruning less informative tokens identified through prediction modules. DVT [51] enhances computational efficiency by automatically adapting the token count for each input image. SuperViT [37] handles diverse patch sizes with a single model, adaptively adjusting token retention during inference.
@@
Existing ViT acceleration methods generally require either fine-tuning all pre-trained parameters [49, 51], or training models from scratch [36, 37]. Consequently, these approaches necessitate substantial training or fine-tuning time to adapt ViT to downstream vision tasks. Recently, Dynamic Tuning (DyT) [60] keeps the pre-trained ViT parameters frozen, updating only the adapters and token dispatchers to enhance parameter efficiency and reduce redundant computation during inference. Unlike DyT, which directly skips the uninformative tokens, our method consolidates these tokens into a single representative token to retain visual features beneficial for classification. Additionally, unlike DyT, which necessitates computing all tokens to update the parameters of the proposed token dispatchers, our approach does not introduce any additional modules for token sparsification. Our method efficiently fine-tunes the pre-trained ViT by selectively adapting the tokens, thereby improving efficiency during both the fine-tuning and inference stages.
@@ 4.2 Main Results
Comparisons on Complete Datasets. We conduct experiments on comprehensive image and video datasets to evaluate the adaptation performance with abundant training data. The results on complete image and video datasets are shown in Table 2, from which we find that: (1) Sparse-Tuning outperforms all baseline methods on both image and video datasets, demonstrating its strong transferability on complete datasets. (2) Sparse-Tuning demonstrates exceptional inference efficiency on both image and video datasets. Particularly on video datasets, Sparse-Tuning reduces the computational complexity of the original ViT-B by around 30%, highlighting its strong efficiency in video applications. With only 1.11M updated parameters, our Sparse-Tuning achieves superior performance in image and video recognition, while significantly improving inference efficiency.
@@ A.1 Performance and Efficiency on CIFAR-100
Full fine-tuning 85.8 18.82 3.43 90 8 17.58 90.91
Adapter [21] 1.19 13.74 2.61 70 8 17.81 91.76
LoRA [22] 1.19 14.07 2.89 75 8 17.58 91.42
AdaptFormer [5] 1.19 13.69 2.61 68 8 17.81 92.03
VPT [25] 0.07 14.60 3.12 77 8 18.32 91.64
DyT [60] 4.80 11.32 2.35 90 7 12.21 91.01
Sparse-Tuning 1.10 8.89 1.53 40 4 11.70 92.31

### v2

@@ Abstract
Parameter-efficient fine-tuning (PEFT) has emerged as a popular solution for adapting pre-trained Vision Transformer (ViT) models to downstream applications. While current PEFT methods have achieved parameter efficiency, they overlook the efficiency of computation and GPU memory during both fine-tuning and inference, falling short of practical requirements. In this paper, we propose Sparse-Tuning, a novel PEFT method that accounts for the information redundancy in images and videos to boost the above efficiency. By sparsely preserving the semantic-relevant tokens and merging irrelevant ones, Sparse-Tuning minimizes the quantity of tokens processed at each layer, leading to a quadratic reduction in computational and memory overhead. To align our token sparsification strategy suitably with fine-tuning purposes, we further design Dense Adapters that establish dense connections from shallow layers to deeper layers. These Dense Adapters integrate multi-level local features to enrich the current tokens, improving both token preservation and model adaptation. Empirical results on VTAB-1K, three image datasets, and two video datasets show that our Sparse-Tuning reduces GFLOPs to 62%-70% of the original ViT-B while achieving state-of-the-art performance. Source code is available at https://github.com/liuting20/Sparse-Tuning.
@@ 1 Introduction
Large-scale Vision Transformer (ViT) models [9, 36, 42, 26] have demonstrated strong generalization capabilities across a wide range of downstream vision tasks. The prevalent approach to adapt these models for specific tasks follows the pretrain-then-finetune paradigm, where models are initially pre-trained on large-scale datasets [7, 33] and then fine-tuned for each downstream task. However, as these pre-trained ViT models continue to scale up [51, 6], fully fine-tuning them becomes more computationally intensive. Additionally, there are risks of catastrophic forgetting and overfitting when fine-tuning on limited downstream datasets [54, 20, 8].
Recently, various parameter-efficient fine-tuning (PEFT) methods [17, 29, 18, 21, 5] have been proposed to address the high computational costs and risks associated with fully fine-tuning large models. By updating additional parameters inserted into the model [17, 5] or appended to the input data [29, 21], PEFT methods can achieve similar or even better performance compared to full fine-tuning. However, we have discovered that the parameter efficiency, pursued by these methods, does not directly correlate with the efficiency that is truly of concern during deployment, including computational and memory efficiency. As shown in Figure 1, despite the significant disparity in the number of trainable parameters, the gap of GFLOPs and memory between PEFT methods and full fine-tuning is not pronounced. The reason lies in the fact that the reduction of trainable parameters achieved by existing PEFT methods does not address the actual bottleneck of ViTs. The computational and memory costs of pre-trained ViTs comes from not only its own size, but also the length of the input due to the quadratic complexity of the attention operation, which is the core component of the Transformer. Thus, we suggest that PEFT methods should consider managing the number of tokens for efficient and effective ViT adaptation.
According to the above analysis, we propose Sparse-Tuning, a novel tuning paradigm that achieves efficient fine-tuning and inference for ViT adaptation. Inspired by the fact that heavy redundancy exhibits spatially in images and spatio-temporally in videos [14, 10], we introduce a parameter-free Token Sparsification method that effectively compresses the extracted visual tokens during fine-tuning. Specifically, it progressively preserves semantic-relevant tokens while condensing irrelevant ones into a representative token, thereby reducing the number of tokens in each layer. While reducing computational overhead, our ablation study (shown in Table 3) indicates that the token reduction may lead to information loss and decreased accuracy. Since the remaining visual tokens in deeper layers can be viewed as low-resolution representations, which contain coarse and holistic characteristics of visual cues, our tuning method further supplements high-resolution local details with well-designed Dense Adapters. Different from conventional intra-layer adapters, our Dense Adapters establish dense connections between different layers, integrating multi-level features from shallower layers to enhance the current tokens. Additionally, the full utilization of local features also contributes to the precise identification of semantic-relevant tokens within the subsequent layers. With these non-trivial designs, Sparse-Tuning improves performance while significantly reducing computational cost and GPU memory consumption for efficient ViT fine-tuning and inference, as shown in Figure 1.
@@
To fully evaluate the generalization, we conduct extensive experiments on the common PEFT benchmark VTAB-1K [52], three complete image datasets: CIFAR-100 [28], SVHN [11], and Food-101 [2], as well as two complete video datasets: Kinetics-400 (K400) [4] and Something-Something V2 (SSv2) [12]. Empirical results on VTAB-1K demonstrate that with only 11.65 GFLOPs, approximately 66% of the computational cost of the original ViT-B, Sparse-Tuning outperforms all state-of-the-art methods in performance and efficiency. Moreover, Sparse-Tuning achieves superior performance in both image and video recognition on complete datasets while significantly improving both fine-tuning and inference efficiency.
@@ 2.1 Parameter-efficient Fine-tuning
With the trend of scaling up Vision Transformers (ViT) [9, 14, 51, 6] for enhanced performance and generalization, adapting the entire model to downstream tasks becomes increasingly computationally expensive. To mitigate this, parameter-efficient fine-tuning (PEFT) [17, 18, 21, 48] emerges as a strategic approach. PEFT methods updates only a small subset of additional parameters while keeping the majority of the pre-trained model frozen, thereby mitigating the risks of catastrophic forgetting and overfitting. Most PEFT methods methods designed for Transformer [44] can be classified into three types: (1) Partially Tuning [27, 50], that only updates a small subset of inherent parameters while freezing most original parameters. (2) Prompt Tuning [29, 21, 20, 47], that integrates a fixed length of learnable tokens (i.e., prompts) appended with the input data. It only updates the prompts during fine-tuning. (3) Adapter Tuning [17, 5, 46, 34, 35], that only updates additional parameters in the module inserted into the model (i.e., Adapter) during fine-tuning.
@@
While most PEFT methods improve parameter efficiency during fine-tuning, they often introduce new parameters that compromise inference efficiency. Reparameterization methods, such as LoRA [18] and FacT [25], introduce learnable parameters that can be integrated into the original model during inference, thus maintaining the original model’s inference efficiency. However, current PEFT methods still fall short in enhancing inference efficiency, which is crucial for adapting large-scale ViTs (e.g., ViT-L) to practical applications. In this paper, we aim to improve both the fine-tuning and inference efficiency of pre-trained ViT.
2.2 Token Compression for ViT
@@
Recent works have explored to accelerate the inference efficiency of ViT [43, 31, 32], with most of them aiming to reduce the token redundancy to decrease computational complexity. For instance, DynamicViT [43] efficiently sparsifies ViT by pruning less informative tokens identified through prediction modules. DVT [45] enhances computational efficiency by automatically adapting the token count for each input image. SuperViT [32] handles diverse patch sizes with a single model, adaptively adjusting token retention during inference.
@@
Existing token compression methods for ViT generally require either fine-tuning all pre-trained parameters [43, 45], or training models from scratch [31, 32]. Consequently, these approaches necessitate substantial training or fine-tuning time to adapt ViT to downstream vision tasks. Recently, Dynamic Tuning (DyT) [55] keeps the pre-trained ViT parameters frozen, updating only the adapters and token dispatchers to enhance parameter efficiency and reduce redundant computation during inference. Unlike DyT, which directly skips the uninformative tokens, our method consolidates these tokens into a single representative token to retain visual features beneficial for classification. Additionally, Different from DyT that necessitates computing all tokens to update the parameters of the proposed token dispatchers, we do not introduce any additional modules for token sparsification. Our method efficiently fine-tunes the pre-trained ViT by selectively adapting the tokens, thereby improving efficiency during both the fine-tuning and inference stages.
@@ 4.2 Main Results
Comparisons on Complete Datasets. We conduct experiments on comprehensive image and video datasets to evaluate the adaptation performance with abundant training data. The results on complete image and video datasets are shown in Table 2, from which we find that: (1) Sparse-Tuning outperforms all baseline methods on both image and video datasets, demonstrating its strong transferability on complete datasets. (2) Sparse-Tuning demonstrates exceptional inference efficiency on both image and video datasets. Particularly on video datasets, Sparse-Tuning reduces the computational complexity of the original ViT-B by around 30%, highlighting its strong efficiency in video applications. With only 1.10M updated parameters, our Sparse-Tuning achieves superior performance in image and video recognition, while significantly improving inference efficiency.
@@ A.1 Performance and Efficiency on CIFAR-100
Full fine-tuning 85.8 17.93 2.57 90 8.03 17.58 90.91
Adapter [17] 1.19 13.74 2.61 75 8.16 17.81 91.76
LoRA [18] 2.16 14.07 2.79 72 8.04 17.58 91.42
AdaptFormer [5] 1.19 13.26 2.61 68 8.09 17.81 92.03
VPT [21] 0.07 14.60 2.80 77 8.24 18.32 91.64
DyT [55] 4.80 14.32 2.15 95 7.12 12.21 91.01
Sparse-Tuning 1.10 8.89 1.53 40 4.13 11.70 92.31